OpenStudy (anonymous):

beth is walking around a circular track she knows she is 10 yards from a tree that is in the middle of the track if beth walks once around the track how far has she walked

OpenStudy (anonymous):

What we are looking for is the circumference of the circle, do you know what the formula of the circumference of a circle is based on the radius?

OpenStudy (anonymous):

2(3.14)r 2(3.14)10 62.8

OpenStudy (anonymous):

Thank you, the formula is c = 2(pi)r

OpenStudy (anonymous):

Okay, so we know r=10 so c=2*pi*10 = 20pi